OpenStudy (anonymous):

Identify the line that contains the error in this solution and correct the work for that line. Line 1: - 2 (x + 1) - 3 is greater than or equal to 11 Line 2: - 2 x + - 2 - 3 is greater than or equal to 11 Line 3: - 2x + - 5 is greater than or equal to 11 Line 4: - 2 x is less than or equal to 16 Line 5: x is less than or equal to negative 8

OpenStudy (mathstudent55):

Can you use numbers and single letters for variables instead of spelling everything out?

OpenStudy (anonymous):

I copied and pasted it that's why it did that @mathstudent55

OpenStudy (anonymous):

im changing it right now

OpenStudy (mathstudent55):

Ok. Thanks.

OpenStudy (anonymous):

I changed it @mathstudent55

OpenStudy (mathstudent55):

Ok, great. Now I can see it. It's much easier to read it this way. Also, now I see there are parentheses in the first line, which did not show before.

OpenStudy (anonymous):

it bc it is solving it so it gets smaller and smaller

OpenStudy (mathstudent55):

Line 1: \(- 2 (x + 1) - 3 \ge 11 \) The first step is to distribute the -2 on the left side. To do that, you multiply -2 by x and multiply -2 by 1. You get: Line 2: \(-2x + (- 2) - 3 \ge 11\) This is exactly what line 2 has above, so line is correct.

OpenStudy (mathstudent55):

Now let's simplify line 2 and see what we get: \(-2x + (- 5) \ge 11\) This is exactly what line 3 above has, so it's still correct.

OpenStudy (mathstudent55):

Line 3 above is correct. Let's now move the -5 to the right side. Since on the left side, there is a -5 being added, which is the same as subtracting -5, we need to do the opposite, so we add 5, but we must do it to both sides of the inequality. Now we add 5 to both sides. \(-2x + (-5) + 5 \ge 11 + 5\) Now we combine the numbers on both sides: \(-2x \ge 16\) Now we notice that this line is different form line 4 above. Here we have a greater than or equal sign, and in line 4 above there is a less than or equal sign. We did our step correctly. The error is above. When you add or subtract a negative number from both sides of an inequality, the inequality sign DOES NOT change. We kept the greater than or equal sign as it was. Above the sign changed from >= to <=. That was an error.

OpenStudy (anonymous):

im doing a essay question so what would I wright

OpenStudy (mathstudent55):

Now we continue: \(-2x \ge 16\) Since x is beiong multiplied by -2, we need to divide both sides by -2 to get x alone. Since we are dividing both sides by a negative number, the direction of the inequality sign changes now. \( \dfrac{-2x}{-2} \le \dfrac{16}{-2} \) Now we do the two divisions and we get the final answer: \(x \le -8\) Line 5 above is correct. There were two errors made above. By coincidence, the two errors "canceled each other" and the final solution to the inequality came out correct.

OpenStudy (anonymous):

so what would I wright

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

You need to write the correct version of line 4: \(-2x \ge 16\) That is the only line given above that is incorrect.

OpenStudy (mathstudent55):

Notice the spelling of "write," "right," and "rite." You can be "right" or wrong. You can go "right" or left. You can "write" and erase. You can celebrate a "rite" of passage. You can't be "wright," or go "wright," or "wright," or celebrate a "wright" of passage. "Wright" is an Old English word that has a different meaning, such a playwright.